OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 14 terms?

OpenStudy (anonymous):

okay, geometric sequences have to be like this: a_2/a_1 = r & a_3/a_2 = r...... and so on

OpenStudy (anonymous):

after you do that, you need to use this equation: S_n = a_1(1-r^n)/(1-r)

OpenStudy (anonymous):

r = 3 so: S_14 = (1- 3^14)(1-3) = -4782968/-2 = 2,391,484

OpenStudy (anonymous):

hope that helps